An empty parallel plate capacitor is connected between the terminals of a 10.6-V battery and charges up. The capacitor is then disconnected from the battery, and the spacing between the capacitor plates is doubled. As a result of this change, what is the new voltage between the plates of the capacitor?
Answer #1

as battery is discharged, Q = constant


Q = CV


C = epsilon * Area / D


if spacing is doubled means, Capicatance becomes half


i.e Voltage becomes doubled


21.2 V is the answer
Answer #2

as battery is discharged, Q = constant


Q = CV


C = epsilon * Area / D


if spacing is doubled means, Capicatance becomes half

C1V1= C2V2

S0 V2 = 2 * V1

i.e Voltage becomes doubled


New voltage = 2 * 10.6 =

21.2 V <------------------------(answer)